I really need help solving this problem by either substitution or elimination, please!
2x=3y-7
3x-4y=-9

Answer by mananth(16949)

2x-3y=-7 .............1
3x-4y=-9 .............2

multiply (1)by -4
Multiply (2) by 3
-8x+12y=28
9x-12y=-27

Add the two equations

x=1
plug value of x in (1)

2x-3y=-7
2-3y=-7
-3y=-7-2
-3y=-9
y=3